OpenStudy (anonymous):

Of the 220 books on a bookshelf, exactly 60 are nonfiction and 110 are hardcover. If 60% of the nonfiction books are hardcover, how many of the books are both not nonfiction and not hardcover? A. 36 B. 50 C. 74 D. 86 E. 110

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Do you know how to answer, because I'm confused.

OpenStudy (anonymous):

Is this a serious question?

jigglypuff314 (jigglypuff314):

yeah, he doesn't know how to do it. I was the one who gave him this to keep him from posting stupid ones.

OpenStudy (anonymous):

I've tried to figure it out, but this is a hard one.

OpenStudy (anonymous):

Hint: apply the 60% to the number of nonfiction books.

OpenStudy (texaschic101):

60% of 60 = .6(60) = 36 (60% of non-fiction are hardcover) 60 non fiction 110 + 36 are hardcover 110 + 36 + 60 = 206 220 - 206 = 14 I am not getting the right answer

OpenStudy (anonymous):

Then try to figure out how many remain when you subtract them out.

OpenStudy (anonymous):

That would be 36 @wio

OpenStudy (texaschic101):

36 + 110 = 146 220 - 146 = 74 ??

jigglypuff314 (jigglypuff314):

not 74

OpenStudy (anonymous):

The number that are hardcover OR nonfiction would be: nonfiction + hardcover - nonfiction and hardcover. You have to eliminate double counting. Then you can finally find out how many are both not hardcover and not nonfiction.

OpenStudy (anonymous):

I still can't figure this out.

OpenStudy (texaschic101):

60 are non fiction...of those 36 are hardcover 110 are hardcover 110 + 60 = 170 220 - 170 = 50 ??

OpenStudy (texaschic101):

the 36 hardcover out of the non-fiction doesn't even matter because they are still non-fiction, right ?

OpenStudy (texaschic101):

I can understand why you got confused

OpenStudy (anonymous):

So, 110+60=170. .60*60=36 170-36=134 220-134=86 Did I get everything, @jigglypuff314?

OpenStudy (anonymous):

Okay 220 total 60 are nonfiction 110 are hardcover 60 * 0.60 = 36 are both nonfiction and hardcover 60 - 36 = 24 are nonfiction but not hardcover 24 + 110 = 134 are either nonfiction or hardcover 220 - 134 = 86 are neither nonfiction nor hardcover

jigglypuff314 (jigglypuff314):

go @wio !!! :D

OpenStudy (anonymous):

Good job @wio.

OpenStudy (texaschic101):

I was way off...my bad

OpenStudy (anonymous):

Can anybody give me a medal, I read it raises your smartscore.

OpenStudy (texaschic101):

done deal :)

OpenStudy (anonymous):

Thanks! That was a tough question.

jigglypuff314 (jigglypuff314):

I've got more if you want @OrangeMaster

OpenStudy (anonymous):

Thanks for the practice @jigglypuff314.

OpenStudy (texaschic101):

I know...I got so confused. Wio makes it look so easy...lol

OpenStudy (anonymous):

Sure I guess one more.